40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15903     Entry time: Thu Mar 11 14:03:02 2021
Author: gautam 
Type: Update 
Category: LSC 
Subject: AO path 

There is some evidence of weird saturation but the gain balancing (0.8dB) and orthogonality (~89 deg) for the daughter board on the REFL11 demod board that generates the AO path error signal seem reasonable. This board would probably benefit from the AD797-->Op27 and thick-film-->thin film swap but i don't think this is to blame for being unable to execute the RF transition.

Attachment 1: IMG_9127.HEIC  2.581 MB
ELOG V3.1.3-